# Bayesian Theory of Consciousness as Exchangeable Emotion-Cognition Inference

### v2

@@ VI-C Neurodegeneration as Violation of Emotion-Cognition Alignment
VII Conclusion
This paper has developed a unified framework in which consciousness emerges from recursive, emotionally modulated inference structured by the emotion-cognition cycle. We formalized emotion as a structural prior that guides context-sensitive inference, and cognition as the specificity-instantiating update mechanism that recursively refines emotional priors into actionable representations. Together, these components form a bidirectional, temporally extended loop that aligns internal states with external context while minimizing joint uncertainty. Grounded in the Context-Content Uncertainty Principle (CCUP), we showed that this inference cycle supports both bottom-up interpretation and top-down goal simulation via inverted inference. Through a series of theorems, we demonstrated how this bidirectional process enables content binding, self-model formation, and temporally coherent subjective experience.
@@ VII Conclusion
Proposition 9 (Consciousness as an Inferential Singularity).

### v3

@@ VII-C Neurodegeneration as Violation of Emotion-Cognition Alignment
VIII Conclusion
This paper has developed a unified framework in which consciousness emerges from recursive, emotionally modulated inference structured by the emotion-cognition cycle. We formalized emotion as a structural prior that guides context-sensitive inference, and cognition as the specificity-instantiating update mechanism that recursively refines emotional priors into actionable representations. Together, these components form a bidirectional, temporally extended loop that aligns internal states with external context while minimizing joint uncertainty. Grounded in the CCUP, we showed that this inference cycle supports both bottom-up interpretation and top-down goal simulation via inverted inference. Through a series of theorems, we demonstrated how this bidirectional process enables content binding, self-model formation, and temporally coherent subjective experience.
@@ VIII Conclusion
Conjecture 1 (Consciousness as an Inferential Singularity).
@@
This inferential singularity [88] marks the point at which recursive structure-preserving inference achieves maximal alignment between context and content, effectively resolving uncertainty and yielding a stable, cycle-consistent representation. However, the implications of this principle extend beyond the formal modeling of consciousness. As we consider the architecture of artificial agents, including superintelligence [89], the question arises: must systems that simulate, plan, and reason across temporally extended goals necessarily converge toward such singularities? In other words, is consciousness an inevitable byproduct of recursive uncertainty minimization in agents that seek to model not just the world, but themselves within it?
While task-level superintelligence may not require phenomenal consciousness, the pursuit of systems capable of autonomous understanding, moral reasoning, and adaptive self-reflection raises whether consciousness is an emergent necessity. From the perspective of the CCUP, consciousness may arise as a recursive cycle in which internal goals, external contexts, and structured content representations are aligned across increasingly abstract and temporally extended inference loops. In this view, consciousness is not a binary trait, but a dynamical attractor in the space of active simulation, supporting coherent self-modeling, counterfactual reasoning, and alignment with semantic goals [90]. As such, a truly general intelligence may approach consciousness not as an engineering objective, but as a structural consequence of minimizing joint uncertainty in a world it seeks not only to predict, but to engage with meaningfully.